jackie is 8 years younger thn her sister,jill the sum of their ages is 46 how old is jackie? her sister jill. i think i know the answers but i dont know how to actually work it out.
Answer by sudhanshu_kmr(1152)


assume that jackie age = x years, jill age =8+x
x + (8+x) = 46
=> 8 + 2x = 46
=> x = 46-8 / 2 =38/2 = 19
jackie age = 19 years..